OpenStudy (anonymous):

For which equation would you subtract 3, and then multiply by 9 to solve? (see reply bar for multiple choice)

OpenStudy (anonymous):

\[A. 9k + 3 = 40 \] \[B. \frac{ k }{ 9 } + 3 = 40\] \[C. 9k - 3 = 40 \] \[D. \frac{ k }{ 9 } - 3 = 40 \]

OpenStudy (anonymous):

So we are trying to find a number which is k. Go from last operation to first operation. If you multiply an expression by 9 and find our value (k) that expression should be k/9. If you substract 3 from another expression and find k/9 it should be k/9+3. The answer is k/9+3 = something (B)

OpenStudy (anonymous):

Thank youuuuu!